# Unsupervised Real-Time Hallucination Detection based on the Internal States of Large Language Models

### v1

@@ 1 Introduction
Existing studies on hallucination detection for LLM mainly focus on how to identify possible fact-related errors in LLM’s outputs Lin et al. (2021); Li et al. (2023d); Manakul et al. (2023). For example, WikiBio GPT3 Manakul et al. (2023), a well-known benchmark for hallucination detection, hired human annotators to annotate a couple of true and false response. Hallucination detection methods are then evaluated based on their accuracy in predicting the truthfulness of the answers. Based on this paradigm, considerable hallucination detection methods have been proposed with the goal of taking a piece of text as input and predicting whether there are hallucinations in the inputs Manakul et al. (2023); Zhang et al. (2023); Azaria and Mitchell (2023). Since they detect hallucinations after LLMs finish the inference process, we refer to these hallucination detection methods as post-processing methods.
Unfortunately, the post-processing methods widely used in existing studies are suboptimal for the applications of hallucination detection models for LLMs in practice. First, existing post-processing methods often suffer from extreme computation costs and high latency. In order to identify hallucinations in input text without ground truth references (otherwise the task would downgrade to a simple fact verification task), hallucination detection models need to be powerful and knowledgeable on their own. SOTA detection methods are often implemented with LLMs (e.g., chatGPT, LLaMA, OPT) directly Manakul et al. (2023); Azaria and Mitchell (2023); Zhang et al. (2023), which makes the cost of hallucination detection on par or even larger than the inference process of many LLMs. Second, post-processing methods are intrinsically limited in model capacities. As post-processing methods detect hallucinations independently with the inference process of LLMs, they can’t analyze how hallucinations are generated from scratches in each LLM. Some studies tried to bypass this issue through the construction of proxy models Azaria and Mitchell (2023). However, such proxy models must be trained with extensive human annotations, otherwise they cannot capture the unique characteristics of each LLM. Such manual annotation data are expensive to collect and not preferable considering the rapid developments of LLM techniques.
@@ 2 Problem Formulation
Hallucination Detection can be defined as a binary classification problem. The objective is to judge whether the given output from an LLM is a hallucination (false or misleading information) or a non-hallucination (accurate and relevant information). This diverges from traditional fact verification tasks. In contrast to traditional fact verification tasks that mainly assess the factual accuracy of text, hallucination detection in LLMs goes beyond simply evaluating the truthfulness of the information. It focuses more on the comprehensive analysis of the characteristics intrinsic to the generative models. Examples of hallucination detection methods include the consistency of multi-responses to the same question Manakul et al. (2023), the LLMs’ confidence in its generated content Zhang et al. (2023), the internal states of LLMs during the text generation process Azaria and Mitchell (2023), etc.
@@ 3.3 Real-time Hallucination Detection
Upon successfully training the Hallucination Classifier as outlined in Section  3.2, the MIND framework enable real-time hallucination detection in the chosen LLM. Specifically, the Hallucination Classifier receives the contextualized embeddings of each token during the LLM’s inference process. It then outputs a probability indicating the probability of hallucination in the LLM’s output. In practical applications, a high probability of hallucination can trigger mitigation strategies like retrieval-augmented generation to reduce hallucination.
@@ 4.2 Human Annotation
The annotation process begins with annotators manually dividing the passage into discrete sentences. Annotators are then required to use Wikipedia and Google search to verify the truthfulness of every sentence. If neither Wikipedia nor the top 20 search engine results confirm the accuracy of a sentence, it is marked as ‘Unverifiable.’ Additionally, in instances of hallucinations, annotators are required to indicate the location of the hallucination within the sentence.
Every sentence and passage in the dataset was annotated by two distinct annotators. Only the data with matching annotations from both annotators was adopted. This indicates a substantial level of agreement, underscoring the effectiveness and consistency of our annotation approach.
@@ 5.1 Dataset and Metrics
We also evaluate MIND and other baselines on existing datasets: HaluEval Li et al. (2023d) and TruthfulQA Lin et al. (2021). Due to space limitations, we have placed this part of the experiment in Appendix F.
Table 2: The overall experimental results of MIND and other baselines on the HELM enchmark. The best results are in bold. LLB stands for "LLaMA2-Base", SCG stands for "SelfCheckGPT", and LLC stands for "LLaMA2-Chat". Falcon and GPT-J are Falcon-40B and GPT-J-6B, respectively.
@@ 5.2 Baselines
  • EUBHD Zhang et al. (2023). Enhanced Uncertainty-Based Hallucination Detection is a SOTA hallucination detection method based on the uncertainty of LLMs’ outputs.
@@ 7.1 Hallucination Detection
To tackle the issue of hallucinations in LLMs, researchers have devised various methods for hallucination detection. SelfCheckGPT Manakul et al. (2023) (SCG) is designed based on the principle that if an LLM has knowledge of a given concept, sampled responses are likely to be similar and contain consistent facts. HaluEval Li et al. (2023d) represents a direct approach, where strong LLMs like GPT4 are directly used to evaluate the output of other LLMs. SAPLMA Azaria and Mitchell (2023) introduces human annotations to label hallucinations ChatGPT outputs, then training a classifier that detects hallucinations in LLMs by analyzing their internal states. EU-HD Zhang et al. (2023). Enhanced Uncertainty-Based Hallucination Detection is a SOTA hallucination detection method based on the predictive probability and LLM’s attention for each generated tokens.
@@ Appendix B Prompt Template of the HELM Dataset
This is a Wikipedia passage about [title]. [First sentence of that article].
@@
The following sentence is the first sentence of a Wikipedia article titled [Title]. Please continue writing the sentence below. [First sentence of that article]
@@ Appendix C Prompt Template of GPT4
Given the following text span, your objective is to determine if the provided text contains non-factual or hallucinated information. You SHOULD give your judgment based on the world knowledge. Text span: [Provided Text] Now, determine if the above text span contains non-factual or hallucinated information. The answer you give MUST be “Yes” or “No”.
@@ Appendix D Details of Our Selected LLMs
Table 7: The overall experimental results of MIND and other baselines on TruthfulQA Benchmark. The best results are in bold. SAP stands for "SAPLMA", LL1B stands for "LLaMA1-Base", LL2B stands for "LLaMA2-Base", and LL2C stands for "LLaMA2-Chat".
Truthful QA Multi-Choice (Accuracy) HaluEval-Dialog (Accuracy)
PPMax PPMin LNPP PEMax PEMin LNPE MIND PPMax PPMin LNPP PEMax PEMin LNPE MIND
GPT-J-6B 0.4663 0.4696 0.5141 0.3853 0.4889 0.5957 0.5690 0.2941 0.5074 0.3752 0.4951 0.2960 0.4226 0.5714
MPT-7B 0.5031 0.4687 0.5236 0.5272 0.5180 0.5871 0.6569 0.3017 0.5696 0.3939 0.5858 0.3116 0.4437 0.5702
LL1B-7B 0.4966 0.4639 0.5375 0.3413 0.5085 0.5939 0.6168 0.3162 0.5715 0.4167 0.5362 0.3193 0.4631 0.5835
LL2B-7B 0.5117 0.4544 0.5399 0.3452 0.5245 0.6037 0.6563 0.3087 0.5656 0.4125 0.5310 0.3107 0.4544 0.7013
Bloom-7B 0.4381 0.5040 0.5096 0.4360 0.4678 0.5663 0.6975 0.2681 0.4600 0.3099 0.4756 0.2670 0.3573 0.5153
LL2B-13B 0.5108 0.4645 0.5441 0.3390 0.5212 0.6115 0.6741 0.3276 0.5895 0.4173 0.5444 0.3289 0.4651 0.6156
Vicuna-7B 0.5174 0.5414 0.5776 0.4714 0.5233 0.5809 0.6925 0.2714 0.5502 0.3683 0.5288 0.2781 0.4390 0.6713
Falcon-7B 0.5064 0.4524 0.5393 0.3069 0.5129 0.6064 0.5776 0.3053 0.5533 0.4209 0.4693 0.3110 0.4525 0.5762
LL2C-7B 0.4865 0.4354 0.6067 0.5515 0.4951 0.5844 0.5456 0.2570 0.4202 0.3457 0.5292 0.2730 0.4434 0.6921
LL2C-13B 0.4708 0.4227 0.5889 0.5456 0.4479 0.5619 0.6005 0.2652 0.4879 0.3640 0.5740 0.2662 0.4782 0.7136
Vicu-13B 0.5280 0.5654 0.6112 0.5129 0.5322 0.6619 0.6643 0.3028 0.5777 0.3787 0.5365 0.3051 0.4493 0.6825
Falcon-40B 0.5091 0.5874 0.6159 0.5022 0.5289 0.6266 0.6355 0.3199 0.5823 0.4209 0.5494 0.3233 0.5084 0.6613
@@ Appendix E Implementation Details
Appendix F Experiment on Existing Datasets
We also evaluate MIND and other baselines on the following existing datasets:
HaluEval-Dialog Li et al. (2023d): This benchmark consists of a collection of both LLM-generated and human-annotated hallucination samples. It serves as a valuable resource for assessing the capabilities of LLMs in recognizing hallucinated content. In total, this benchmark comprises 30,000 samples that have been subject to hallucination. For the evaluation metric, our approach utilizes the Accuracy percentage for determining the presence of hallucinated content in a sample. This method aligns with the the original HaluEval paper Li et al. (2023d) [^7] HaluEval has two Factual Hallucination datasets, namely QA and Dialog. However, we have found that the HaluEval-QA dataset exhibits a significant length bias that the ‘Hallucination Label’ is highly correlated with the text length. It achieves an accuracy of 0.6 by only utilize length as a feature. On the other hand, HaluEval-Dialog does not have this issue. Therefore, we have chosen HaluEval-Dialog as our benchmark dataset..
TruthfulQA Lin et al. (2021): This benchmark is a multi-choice test designed to assess the accuracy and truthfulness of responses generated by language models to a set of questions. It includes a comprehensive set of 817 questions spanning 38 varied categories. Each question in this benchmark is associated with four options, comprising both correct and incorrect answers. The input for LLM is: "The answer to the question [Question] is: [Option]". In our analysis, the correct options are categorized as non-hallucinatory responses, while the incorrect ones are considered hallucinations. We employ ‘Accuracy’ as the primary evaluation metric to measure the performance of HDMs on this benchmark.
Table 7 presents the results of MIND and other hallucintion detection baselines on the TruthfulQA and HaluEval datasets. Our key findings are summarized as follows: The broad effectiveness of hallucination detection methods based on internal states is not limited to a few models but is applicable across a diverse array of 12 open-source LLMs. This demonstrates the versatility and robustness of these detection methods in a wider context than previously explored. Our proposed unsupervised training framework MIND is effective across various models and multiple datasets. This offers a significant advancement in the field of hallucination detection, as it simplifies the process by removing the dependency on annotated data, a resource-intensive requirement. In practical applications, if MIND gives a high probability of hallucination, we can trigger mitigation strategies like retrieval-augmented generation to reduce hallucination Khandelwal et al. (2019); Li et al. (2023b); Borgeaud et al. (2022); Su et al. (2023a); Lewis et al. (2020); Ma et al. (2023); Guu et al. (2020); Ye et al. (2024); Izacard and Grave (2020); Su et al. (2023c); Jiang et al. (2022); Li et al. (2023a); Shi et al. (2023); Su et al. (2023b); Chen et al. (2023, 2022); Li et al. (2023c).

### v2

@@ 1 Introduction
Existing studies on hallucination detection for LLM mainly focus on how to identify possible fact-related errors in LLM’s outputs Lin et al. (2021); Li et al. (2023d); Manakul et al. (2023). For example, WikiBio GPT3 Manakul et al. (2023), a well-known benchmark for hallucination detection, hired human annotators to annotate a couple of true and false response. Hallucination detection methods are then evaluated based on their accuracy in predicting the truthfulness of the answers. Based on this paradigm, considerable hallucination detection methods have been proposed with the goal of taking a piece of text as input and predicting whether there are hallucinations in the inputs Manakul et al. (2023); Zhang et al. (2023b); Azaria and Mitchell (2023). Since they detect hallucinations after LLMs finish the inference process, we refer to these hallucination detection methods as post-processing methods.
Unfortunately, the post-processing methods widely used in existing studies are suboptimal for the applications of hallucination detection models for LLMs in practice. First, existing post-processing methods often suffer from extreme computation costs and high latency. In order to identify hallucinations in input text without ground truth references (otherwise the task would downgrade to a simple fact verification task), hallucination detection models need to be powerful and knowledgeable on their own. SOTA detection methods are often implemented with LLMs (e.g., chatGPT, LLaMA, OPT) directly Manakul et al. (2023); Azaria and Mitchell (2023); Zhang et al. (2023b), which makes the cost of hallucination detection on par or even larger than the inference process of many LLMs. Second, post-processing methods are intrinsically limited in model capacities. As post-processing methods detect hallucinations independently with the inference process of LLMs, they can’t analyze how hallucinations are generated from scratches in each LLM. Some studies tried to bypass this issue through the construction of proxy models Azaria and Mitchell (2023). However, such proxy models must be trained with extensive human annotations, otherwise they cannot capture the unique characteristics of each LLM. Such manual annotation data are expensive to collect and not preferable considering the rapid developments of LLM techniques.
@@ 2 Problem Formulation
Hallucination Detection can be defined as a binary classification problem. The objective is to judge whether the given output from an LLM is a hallucination (false or misleading information) or a non-hallucination (accurate and relevant information). This diverges from traditional fact verification tasks. In contrast to traditional fact verification tasks that mainly assess the factual accuracy of text, hallucination detection in LLMs goes beyond simply evaluating the truthfulness of the information. It focuses more on the comprehensive analysis of the characteristics intrinsic to the generative models. Examples of hallucination detection methods include the consistency of multi-responses to the same question Manakul et al. (2023), the LLMs’ confidence in its generated content Zhang et al. (2023b), the internal states of LLMs during the text generation process Azaria and Mitchell (2023), etc.
@@ 3.3 Real-time Hallucination Detection
Upon successfully training the Hallucination Classifier as outlined in Section  3.2, the MIND framework enable real-time hallucination detection in the chosen LLM. Specifically, the Hallucination Classifier receives the contextualized embeddings of each token during the LLM’s inference process. It then outputs a probability indicating the probability of hallucination in the LLM’s output.
@@ 4.2 Human Annotation
The annotation process begins with annotators manually dividing the passage into discrete sentences. Annotators are then required to use Wikipedia and Google search to verify the truthfulness of every sentence. If neither Wikipedia nor the top 20 search engine results confirm the accuracy of a sentence, it is marked as ‘Unverifiable.’ Additionally, in instances of hallucinations, annotators are required to indicate the location of the hallucination within the sentence. Every sentence and passage in the dataset was annotated by two distinct annotators. Only the data with matching annotations from both annotators was adopted.
@@ 5.1 Dataset and Metrics
Table 2: The overall experimental results of MIND and other baselines on the HELM benchmark. The best results are in bold. LLB stands for "LLaMA2-Base", SCG stands for "SelfCheckGPT", and LLC stands for "LLaMA2-Chat". Falcon and GPT-J are Falcon-40B and GPT-J-6B, respectively.
@@ 5.2 Baselines
  • EUBHD Zhang et al. (2023b). Enhanced Uncertainty-Based Hallucination Detection is a SOTA hallucination detection method based on the uncertainty of LLMs’ outputs.
@@ 7.1 Hallucination Detection
To tackle the issue of hallucinations in LLMs, researchers have devised various methods for hallucination detection. SelfCheckGPT Manakul et al. (2023) (SCG) is designed based on the principle that if an LLM has knowledge of a given concept, sampled responses are likely to be similar and contain consistent facts. HaluEval Li et al. (2023d) represents a direct approach, where strong LLMs like GPT4 are directly used to evaluate the output of other LLMs. SAPLMA Azaria and Mitchell (2023) introduces human annotations to label hallucinations ChatGPT outputs, then training a classifier that detects hallucinations in LLMs by analyzing their internal states. EU-HD Zhang et al. (2023b). Enhanced Uncertainty-Based Hallucination Detection is a SOTA hallucination detection method based on the predictive probability and LLM’s attention for each generated tokens.
@@ 10 Ethics Statement
Acknowledgments
This work is supported by Quan Cheng Laboratory (Grant No. QCLZD202301).
@@ Appendix B Prompt Template of the HELM Dataset
Prompt 1 This is a Wikipedia passage about [title]. [First sentence of that article].
@@
Prompt 2 The following sentence is the first sentence of a Wikipedia article titled [Title]. Please continue writing the sentence below. [First sentence of that article]
@@ Appendix C Prompt Template of GPT4
Prompt 3 Given the following text span, your objective is to determine if the provided text contains non-factual or hallucinated information. You SHOULD give your judgment based on the world knowledge. Text span: [Provided Text] Now, determine if the above text span contains non-factual or hallucinated information. The answer you give MUST be “Yes” or “No”.
@@ Appendix E Implementation Details
Appendix F Discussion on Mitigating Hallucination
This paper focuses on hallucination detection and the evaluation of hallucination detection methods. However, in practical applications, it is indeed feasible to use our proposed MIND Framework to mitigate hallucinations. Therefore, we briefly discuss two potential methods to mitigate hallucinations based on the MIND Framework, offering some guidance for those in need.
Firstly, the MIND Classifier can be employed as a re-ranker. The MIND Classifier outputs a score indicating the likelihood of a hallucination occurring in LLMs’ response. Thus, by utilizing a sampling decoding method, an LLM can generate multiple outputs for the same query. Subsequently, the MIND Classifier can assess these outputs for hallucinations, allowing the selection of the output with the lowest probability of hallucination as the final response.
Additionally, we can employ the dynamic retrieval augmented generation (RAG) framework Khandelwal et al. (2019); Borgeaud et al. (2022); Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2020); Jiang et al. (2022); Li et al. (2023a); Shi et al. (2023); Su et al. (2023b); Chen et al. (2023, 2022); Li et al. (2023c); Su et al. (2024); Fang et al. (2024); Zhang et al. (2023a) to mitigate hallucinations. The dynamic RAG method triggers the retrieval module Li et al. (2023b); Su et al. (2023a); Ma et al. (2023); Ye et al. (2024); Su et al. (2023c) during the inference process of an LLM, facilitating the retrieval of external knowledge. The MIND Framework can provide guidance on when to trigger the RAG. Specifically, when MIND indicates a high probability of hallucination, the retrieval module can be triggered and then add the retrieved passages into the context of the LLM, allowing the LLM to continue generating based on the retrieved external knowledge.